OpenStudy (anonymous):

how do you solve one step inequalities? PLS help will fan and medal best relpy

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

what is a "one step" inequality? do you have an example?

OpenStudy (anonymous):

Solve this as you would a normal equation x + 4 greater than or equal to 10

OpenStudy (anonymous):

then subtract the 4 from both side and you will have your answer

OpenStudy (anonymous):

so x + 4 - 4 greater than or equal to 10 - 4 equals x greater than or equal to 6

OpenStudy (anonymous):

\[X \ge 6\]

OpenStudy (anonymous):

\[X + 4 \ge 10\]

OpenStudy (anonymous):

\[x + 4 - 4\ge 10 - 4\]

OpenStudy (anonymous):

\[x \ge 6\]

OpenStudy (anonymous):

thx